The Green Machine Manufacturing Company has the option to make or buy a component part for one of its lawnmowers. The annual requirement is 25,000 units. A supplier is able to supply the parts for $10 per piece. Green Machine estimates that it will cost $1000 to prepare the contract with the supplier. To make the parts in-house, Green Machine must invest $100,000 in capital equipment. They estimate it will cost $8 per piece to produce the part in-house. Carry all calculations out to two decimal places.

4. What is the breakeven quantity?

A) Less than or equal to 30,000

B) Greater than 30,000 but less than or equal to 45,000

C) Greater than 45000 but less than or equal to 50,000

D) Greater than 50,000 but less than or equal to 65,000 E) Greater than 65,000
Answer #1

C) Greater than 45000 but less than or equal to 50,000

The break even point be X, then

Expense making = Expense while buying

Fixed Cost + Variable cost = Fixed cost + Variable cost

100000 + 8*X = 1000 + 10*X

2*X = 99000

X = 49500